OpenStudy (ally_teddy):

Need help with this trigonometry question! I need someone to show me step by step solutions so i can understand.

OpenStudy (ally_teddy):

What is the exact value of \[\cos \frac{ \pi }{ 6 }\] as found on the unit circle?

OpenStudy (karim728):